OpenStudy (anonymous):

Solve this exponential function. Find x. 16^(4x) ·(1/4)^(8x + 4)= 64 · (4^x)^−1

OpenStudy (binary3i):

x=-1

OpenStudy (binary3i):

you log to solve the equation

OpenStudy (anonymous):

-1 does not work

OpenStudy (binary3i):

sorry its 7

OpenStudy (anonymous):

ok so i just have to plug those numbers into the calculator to get the answer

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

\[16^{4x}*(\frac{ 1 }{ 4 })^{8x+4}=64*(4^{x})^{-1}\]

OpenStudy (anonymous):

Oh its been solved I had to clean it up to see it better.